# State Combinatorial Generalization In Decision Making With Conditional Diffusion Models

### v1

@@ Abstract
Many real-world decision-making problems are combinatorial in nature, where states (e.g., surrounding traffic of a self-driving car) can be seen as a combination of basic elements (e.g., pedestrians, trees, and other cars). Due to combinatorial complexity, observing all combinations of basic elements in the training set is infeasible, which leads to an essential yet understudied problem of zero-shot generalization to states that are unseen combinations of previously seen elements. In this work, we first formalize this problem and then demonstrate how existing value-based reinforcement learning (RL) algorithms struggle due to unreliable value predictions in unseen states. We argue that this problem cannot be addressed with exploration alone, but requires more expressive and generalizable models. We demonstrate that behavior cloning with a conditioned diffusion model trained on expert trajectory generalizes better to states formed by new combinations of seen elements than traditional RL methods. Through experiments in maze, driving, and multiagent environments, we show that conditioned diffusion models outperform traditional RL techniques and highlight the broad applicability of our problem formulation.
@@ 1 Introduction
To facilitate this study, we first provide formal definitions of state combination and OOC generalization. We then demonstrate the challenge of this task by showing how traditional RL algorithms struggle to generalize in this setting due to unreliable value prediction, and the need for a more expressive policy. On the hunt for a suitable solution, we draw inspiration from the linear manifold hypothesis in diffusion models (Chung et al., 2023; He et al., 2024b) and recent advances in combinatorial image generation (Okawa et al., 2024), and present diffusion models as a promising direction by showing how they can naturally account for the combinatorial structure of states into the diffusion process, enabling better generalization in OOC settings.
@@ 2.1 Generalization in RL
The problem of zero-shot domain transfer assumes that the model is trained and tested on different domains that might have some similarities but are sampled from different underlying distributions (Kirk et al., 2023). One widely used technique is domain randomization, approaching this problem by producing a wide range of contexts in simulation (Kirk et al., 2023; Mehta et al., 2020). Although the focus is also unsupported state space, it commonly assumes that information about the testing environment is not accessible (Mehta et al., 2020) and focuses more on sim2real problems (Kirk et al., 2023). Whereas we assume test time information is given through conditioning but restricting the training set to have narrow coverage.
@@ 2.2 Combinatorial Generalization
Computer Vision
The closest line of work to ours is combinatorial generalization for image generation where the model needs to learn new combinations of a discrete set of basic concepts like color and shapes and generalize to unseen combinations (Wiedemer et al., 2024; Okawa et al., 2024; Schott et al., 2021; Hwang et al., 2023). This problem is often approached with disentangled representation learning (Liu et al., 2023; Schott et al., 2021) with models like VAE but little evidence shows they can fully exhibit generalization ability (Schott et al., 2021; Montero et al., 2020). Okawa et al. (2024) studied the capabilities of conditioned diffusion models on a synthetic shape generation task and showed that their composition ability emerges with enough training, first to closer concepts, then to farther ones.
RL
@@ 3.1 States Formed by Element Combinations
Following Wiedemer et al. (2024), we first denote $e\in{\mathbf{E}}$ to be a base element for an environment. A base element is defined to be the most elementary and identifiable element that is relevant to the decision making task of interest. For example, in a traffic environment, the set ${\mathbf{E}}$ can be the set of vehicles that can occur in the environment such as {car, bike}; and in a 2D maze environment, the set ${\mathbf{E}}$ can be the set of possible locations labeled by the $x,y$-axis coordinate of the agent, i.e. $\mathbb{R}^{2}$. Suppose there are a finite number of $n$ base elements in an environment. Since these elements are the fundamental components relevant to the decision making task, we can form a latent vector ${\bm{z}}=(z_{1},z_{2},...,z_{n})\in{\mathbf{Z}}\equiv{\mathbf{E}}^{n}$, where $z_{i}\in{\mathbf{E}}$ $\forall i\in\{1,\dots,n\}$ that represents the combination of all rudimentary components appearing in this environment related to the decision making task.
Each element can also be associated with a collection of attributes ${\bm{r}}$ such as the color of the vehicle and the velocity of the agent. Attributes are components that are necessary for rendering the states and the rendering function $f({\bm{z}},({\bm{r}}_{1},{\bm{r}}_{2},\dots,{\bm{r}}_{n}))$ can then map the latent and the attributes to a state ${\bm{s}}\in{\mathbf{S}}$. In the traffic environment example, $f$ is equivalent to reconstructing the cars and the bikes given their colors and positions, etc. All reconstructed base elements collectively determine a state ${\bm{s}}$. Concretely, we provide the following definition:
@@ 3.2 Generalization on Probability Space Support
[…]
@@ 3.3 Constraint for OOC Generalization
One crucial assumption made by OOC generalization is that all base elements are seen at training time. Recall $\bm{z}=\big{(}z_{1},z_{2},...,z_{n}\big{)}\text{ where }z_{i}\in{\mathbf{E}}\text{ for }i=1,...,n$. This indicates that the marginal distribution $p(z_{i})>0$ for all $z_{i}$ at training time, or equivalently the training probability space has full support over the marginals. For discrete latent spaces, this also implies that every base element that appeared in the sample space would appear at least once in one latent feature $\bm{z}$. To ensure full support of base elements, the union of marginal supports at test time should be a subset of that at training time. Finally, to test generalizability, we assume $\operatorname{supp}\{P_{train}\}\subsetneqq{\mathbf{Z}}$, i.e. the training probability space doesn’t have full support on the entire latent space.
@@ 4 Why traditional RL fails
4.1 RL and expected reward estimaiton
Figure 3: Expected Q value of CQL and actual return-to-go (RTG) in unsupported states in Roundabout environment.
Most deep RL algorithms rely on learning a Q or Value function, which takes in the current state as network input and predicts the expected future reward (Schulman et al., 2017; Haarnoja et al., 2018). Since states with unseen composition are unsupported and fall within the undertrained regions of the neural network, the value prediction is highly unreliable. This affects both value-based methods that directly choose the maximum action with erroneous Q value and policy-based methods that update the actor with an erroneous value prediction. We plot the expected Q-values learned by CQL alongside the actual return-to-go in both failed and success scenarios in Roundabout environment (Leurent, 2018) (Section 7.1) when presented with OOC states in Figure 3. The grey dashed line is the expected Q-values the model predicts for in-distribution states.
One key observation can be made: Q function shows signs of memorizing, which assigns similar Q values for both training and OOC states.
Despite the problem of distribution shift being a central challenge for offline RL (Levine et al., 2020), online methods also suffer from unseen states when zero-shot generalizing to unsupported states. Traditionally distribution shifts are mitigated with a wider training state distribution under the assumption that test time states are sampled from a distribution with different probability density but same support. However, since new states with different object combinations are out of support of the training environment, using a more exploratory online policy or collecting more training trajectories for offline RL will not fundamentally solve this issue. We need a policy with better generalization to unsupported states to achieve zero-shot generalization in this problem.
@@ 5 Why Diffusion Models Generalize Better
We first introduce diffusion model notations and then provide a proof sketch and experimental evidence of why diffusion models can generalize to OOC states.
@@ 5.2 OOC Generalization in Diffusion Models
We demonstrate that a well-trained diffusion model can naturally sample OOC states that satisfy combinatorial support constraint 3.4 with non-zero probability at test time. The key idea is that pseudo-random denoising trajectories can be constructed at inference time that yield OOC samples with non-zero probability.
@@
Since our states are formed by combinations of base elements (Definition 3.1), with well constructed ${\mathbf{Z}}$ we can assume that the states lie on a lower dimensional manifold $\mathcal{M}$ (representing combinations of base elements) embedded in the high dimensional ambient state space. In some cases such as maze navigation where the latent space is a linear subspace, we can even assume that the underlying manifold $\mathcal{M}$ is a linear manifold whose tangent space is isomorphic to itself. With these assumptions, we present the following corollary whose proof is shown in the appendix B.
Corollary 5.1.
Suppose the states lie along a linear manifold $\mathcal{M}$ in the state space ${\mathbf{S}}$ and the latent space ${\mathbf{Z}}$ is well constructed so that ${\mathbf{Z}}$ is (affine) isomorphic to $\mathcal{M}$. Then a diffusion model $p_{\theta}$ that is well trained on $P_{train}$ can sample an OOC state with non-zero probability.
While the linear manifold assumption may not hold for more complex states, recent computer vision research provides evidence of the combinatorial generalization capabilities of diffusion models in more complicated data spaces: Okawa et al. (2024) showed that given different concepts like shape, color, and size in synthetic shape generation, conditional diffusion models demonstrate a multiplicative emergence of combinatorial abilities where it will first learn how to generalize to concepts closer to the training samples (i.e. only change one of color, shape, and size) and eventually adopt full compositional generalization ability with enough training. Aithal et al. (2024) identifies the phenomena where diffusion models generate samples out of the support of training distribution through interpolating different complex modes on a data manifold. Kadkhodaie et al. (2023) demonstrate generalization to unsupported data by showing two diffusion models trained on large enough non-overlapping data converge to the same denoising function. In the next sections, we discuss how to use diffusion models to handle this challenging problem and also provide empirical evidence in decision-making tasks.
@@ 6 Conditioned Planning with Diffusion
6.1 Diffusion for Imaginary Trajectory Prediction
We follow the same setup as in Janner et al. (2022) and learn a conditional diffusion model that denoises (predicts) the future state-action pairs given the current state and the latent vector. This formulation excels at OOC state generalization because the capability of the diffusion model to sample OOC states enables it to generate a better world model prediction. Since future states of an OOC state are almost always going to be another OOC state (same base element but different attributes), the diffusion model is capable of generating reasonable future predictions and thus facilitating planning. This advantage is visually demonstrated in Figure 6.
6.2 Experiment Setup (Methodology / 7.1)
Since each state corresponds to a latent vector, we will use the latent vector, which is oftentimes the category of the base elements in the current state, as conditioning for diffusion models. This information is extracted from observations and is equally accessible to all models. We let the correspondence between observation, action, and latent vector (rendering function $f$) be implicitly learned by the model and incorporate a cross-attention layer after each residual block for Diffusion Unet (Janner et al., 2022) to facilitate learning. The expert offline data is collected for behavior cloning by fully trained PPO agents on each combination of the training environment and includes only successful rollouts. Since expert trajectories are used for training, we model actions together with states under the assumption that unseen states generated from new conditioning will correspond to statistically reasonable actions. The first action in the generated trajectory is then used to step the environment (Appendix D.5). Detailed architecture of the model can be found in Appendix D.3 and the planning process is described in Algorithm 1.

### v2

@@ Abstract
Many real-world decision-making problems are combinatorial in nature, where states (e.g., surrounding traffic of a self-driving car) can be seen as a combination of basic elements (e.g., pedestrians, trees, and other cars). Due to combinatorial complexity, observing all combinations of basic elements in the training set is infeasible, which leads to an essential yet understudied problem of zero-shot generalization to states that are unseen combinations of previously seen elements. In this work, we first formalize this problem and then demonstrate how existing value-based reinforcement learning (RL) algorithms struggle due to unreliable value predictions in unseen states. We argue that this problem cannot be addressed with exploration alone, but requires more expressive and generalizable models. We demonstrate that behavior cloning with a conditioned diffusion model trained on successful trajectory generalizes better to states formed by new combinations of seen elements than traditional RL methods. Through experiments in maze, driving, and multiagent environments, we show that conditioned diffusion models outperform traditional RL techniques and highlight the broad applicability of our problem formulation.
@@ 1 Introduction
To facilitate this study, we first provide formal definitions of state combination and OOC generalization. We then demonstrate the challenge of this task by showing how traditional RL algorithms struggle to generalize in this setting due to unreliable value prediction, and the need for a more expressive policy. On the hunt for a suitable solution, we draw inspiration from the linear manifold hypothesis in diffusion models (Chung et al., 2023; He et al., 2024b) and recent advances in combinatorial image generation (Okawa et al., 2024), and present conditional diffusion models as a promising direction by showing how they can naturally account for the combinatorial structure of states into the diffusion process, enabling better generalization in OOC settings.
@@ 2.1 Generalization in RL
Meta RL
Unsupervised RL
Unsupervised RL aims to acquire broadly useful skills without access to task-specific rewards and then leverage these skills to accelerate or improve performance on downstream tasks with extrinsic rewards, often via an initial unsupervised “exploration” phase followed by a reward-supervised fine-tuning phase (Eysenbach et al., 2018; Laskin et al., 2022; Jaderberg et al., 2016). While our setting also assumes reuse of previously learned information and focuses on generalization, it differs in that we operate under a fully offline, reward-agnostic setup, without access to or reliance on any implicit reward signals during training.
@@
The problem of zero-shot domain transfer assumes that the model is trained and tested on different domains that might have some similarities but are sampled from different underlying distributions (Kirk et al., 2023). The most widely used technique is domain randomization, approaching this problem by producing a wide range of contexts in simulation (Kirk et al., 2023; Mehta et al., 2020). However, it commonly assumes that information about the testing environment is not accessible (Mehta et al., 2020) and focuses more on sim2real problems (Kirk et al., 2023), whereas we assume test time information is given through conditioning but restricting the training set to have narrow coverage.
@@ 2.2 Combinatorial Generalization
In Computer Vision
The closest line of work to ours is combinatorial generalization for image generation where the model needs to learn new combinations of a discrete set of basic concepts like color and shapes and generalize to unseen combinations (Wiedemer et al., 2024; Okawa et al., 2024; Schott et al., 2021; Hwang et al., 2023). This problem is often approached with disentangled representation learning (Liu et al., 2023; Schott et al., 2021) with models like VAE but little evidence shows they can fully exhibit generalization ability (Schott et al., 2021; Montero et al., 2020). Okawa et al. (2024) studied the capabilities of conditioned diffusion models on a synthetic shape generation task and showed that their composition ability emerges with enough training, first to closer concepts, then to farther ones. However, we would like to emphasize the difference between image generation and decision-making tasks, where decision-making in the classic RL setting includes evaluating candidate actions, interacting with unpredictable environments, and handling future state distribution shifts caused by previous action choices and environment (Levine et al., 2020), different from image generation where it is one step and does not interact with the environment.
In RL
@@ 3.1 States Formed by Element Combinations
Following Wiedemer et al. (2024), we first denote $e\in{\mathbf{E}}$ to be a base element for an environment. A base element is defined to be the most elementary and identifiable element that is relevant to the decision making task of interest. For example, in a traffic environment, the set $\mathbf{E}$} can be the set of vehicles that can occur in the environment such as {car, bike}; and in a 2D maze environment, the set $\mathbf{E}$} can be the set of possible locations labeled by the $x,y$-axis coordinate of the agent, i.e. $\mathbb{R}^{2}$. Suppose there are $n$ base elements in an environment. Since these elements are the fundamental components relevant to the decision making task, we can form a latent vector ${\bm{z}}=(z_{1},z_{2},...,z_{n})\in{\mathbf{Z}}\equiv{\mathbf{E}}^{n}$, where $z_{i}\in{\mathbf{E}}$ $\forall i\in\{1,\dots,n\}$ that represents the combination of all rudimentary components appearing in this environment related to the decision making task.
Each element can also be associated with a collection of attributes $\bm{r}$} such as the color of the vehicle and the velocity of the agent. Attributes are components that are necessary for rendering the states but not essential for the decision making task. The rendering function $f({\bm{z}},({\bm{r}}_{1},{\bm{r}}_{2},\dots,{\bm{r}}_{n}))$ then maps the latent and the attributes to a state ${\bm{s}}\in{\mathbf{S}}$. In the traffic environment example, $f$ is equivalent to reconstructing the cars and the bikes given their colors and positions, etc. All reconstructed base elements collectively determine a state $\bm{s}$}. Concretely, we provide the following definition:
@@ 3.2 Generalization on Probability Space Support
Define the sample space $\mathbf{Z}$} as the set of all possible $\bm{z}$. $\sigma$-algebra $\Sigma=\mathcal{B}({\mathbf{Z}})$ is the Borel set of $\mathbf{Z}$}. $p:{\mathbf{Z}}\rightarrow[0,1]$ such that $\int_{{\bm{z}}\in{\mathbf{Z}}}p({\bm{z}})d{\bm{z}}=1$ is the probability dense function (PDF). Then the probability space over the latent vector $\mathbf{z}$ can be defined as $P=({\mathbf{Z}},\Sigma,p)$.
@@ 3.3 Constraint for OOC Generalization
One crucial assumption made by OOC generalization is that all base elements are seen at training time. Recall the latent vector $\bm{z}=\big(z_{1},z_{2},...,z_{n}\big)\text{ where }z_{i}\in{\mathbf{E}}\text{ for }i=1,...,n$ represent the appearing base elements for a given state. This indicates that the marginal distribution $p(z_{i})>0$ for all $z_{i}$ at training time, or equivalently the training probability space has full support over the marginals. For discrete latent spaces, this also implies that every base element that appeared in the training state space would appear at least once in one latent feature $\bm{z}$. To ensure full support of base elements, the union of marginal supports at test time should be a subset of that at training time. Finally, to test generalizability, we assume $\operatorname{supp}\{P_{train}\}\subsetneqq{\mathbf{Z}}$, i.e. the training probability space doesn’t have full support on the entire latent space.
@@ 4 Why traditional RL fails
4.1 RL and expected reward estimation
Most deep RL algorithms rely on learning a Q or Value function, which takes in the current state as network input and predicts the expected future reward (Schulman et al., 2017; Haarnoja et al., 2018). Since states with unseen composition are unsupported and fall within the undertrained regions of the neural network, the value prediction is highly unreliable. This affects both value-based methods that directly choose the maximum action with erroneous Q value and policy-based methods that update the actor with an erroneous value prediction. We plot the expected Q-values learned by CQL alongside the actual return-to-go in both failed and success scenarios in Roundabout environment (Leurent, 2018) (Section 7.1) when presented with OOC states in Figure 3. The grey dashed line is the expected Q-values the model predicts for in-distribution (ID) states. When comparing the Q-value predicted on in-distribution and OOC states:
Figure 3: Expected Q value predicted by CQL and in OOC and ID states actual return-to-go (RTG) in OOC states in Roundabout environment.
One observation about inaccuracies of value prediction can be made: Q function shows signs of memorizing, which assigns similar Q values for both ID and OOC states.
The problem of distribution shift is even more challenging for offline RL, where suboptimal action prediction can cause future states to deviate away from the training set and lead to compounding errors (Levine et al., 2020). This problem can often be mitigated with a wider training state distribution under the assumption that test time states are sampled from a distribution with different probability density but same support. However, since new states with different object combinations are out of support of the training environment, neither using a more exploratory policy nor collecting more training trajectories for offline RL can reach these OOC states, and thus the issue cannot be fundamentally resolved. We need a policy with better generalization to unsupported states to achieve zero-shot generalization in this problem.
@@ 5 Why Diffusion Models Generalize Better
In this section, we first introduce diffusion model notations and then provide a proof sketch of why diffusion models can generalize to OOC states.
@@ 5.2 OOC Generalization in Diffusion Models
We now give a theoretical justification for OOC generalization by showing that a well-trained diffusion model concentrates probability mass on in-combination states (seen and OOC), with a quantifiable density lower bound.
@@
Since states are formed by combinations of base elements (Definition 3.1), with a well-constructed and encoded $\mathbf{Z}$}, we assume that the set of in-combination states (both supported and OOC states) concentrates on a lower-dimensional manifold $\mathcal{M}$ embedded in the high dimensional ambient state space. As a result, we first adopt a linear manifold assumption (Assumption B.1), where $\mathcal{M}$ is the linear manifold that the in-combination states lie along.
To connect this formulation to the diffusion sampling process, we assume we have access to a well-trained diffusion model (Assumption B.2) that: (i) contract the components that are orthogonal to the manifold and pull the samples towards the manifold (block-wise bi-Lipschitz with contraction), and (ii) the learned reserve mean should not mix the manifold orthogonal and on-manifold components (block preserving denoising). With above assumptions, we present the following theorem that quantifies the lower bound of the probability density assigned to a state:
Theorem 5.1.
Under the linear manifold assumption (Assumption B.1), and let $p_{\theta}$ be a DDPM diffusion using parametrization from Equation 2 that is well trained (Assumption B.2), with some constant $C$, for all ${\bm{s}}\in S$
$p_{\theta}({\bm{s}})\geq C\exp{\left(-\frac{\|W_{\mathcal{M}}({\bm{s}}-\mu_{0})\|^{2}}{2(\sigma_{0}^{\mathcal{M}})^{2}}-\frac{\|W_{\perp}{\bm{s}}\|^{2}}{2(\sigma_{0}^{\perp})^{2}}\right)}$ (1)
where $0<A_{t}^{\mathcal{M}}\leq B_{t}^{\mathcal{M}}<\infty$, $0<A_{t}^{\perp}\leq B_{t}^{\perp}<1$ are the block-wise bi-Lipschitz constants for the on-manifold component and the orthogonal component respectively, $W_{\mathcal{M}}$ and $W_{\perp}$ are projection operators onto the manifold and its orthogonal complement respectively, $\sigma_{0}^{\mathcal{M}}$ and $\sigma_{0}^{\perp}$ are the effective variances along and orthogonal to the manifold, $\mu_{0}$ is the learned noise-free center, and $C=(2\pi)^{-d/2}\left(\prod_{t=1}^{T}(B_{t}^{\mathcal{M}}/A_{t}^{\mathcal{M}})^{-m}(B_{t}^{\perp}/A_{t}^{\perp})^{m-d}\right)(\sigma_{0}^{\mathcal{M}})^{-m}(\sigma_{0}^{\perp})^{m-d}$.
Proof sketch: The proof proceeds by induction over diffusion timesteps $t=T,\ldots,0$. We begin with the base case at $t=T$ where $\mathbf{s}_{T}\sim\mathcal{N}(0,I)$, and establish the initial Gaussian density lower bound that factorizes along the manifold and its orthogonal complement via the projectors $(W_{\mathcal{M}},W_{\perp})$. For the inductive step from timestep $k$ to k1k-1, we decompose the DDPM reverse process $\mathbf{s}_{k-1}=\bm{\mu}_{\theta}(\mathbf{s}_{k},k)+\sigma_{k}\epsilon$ into two operations: (1) applying the learned denoising mean $\bm{\mu}_{\theta}(\cdot,k)$, and (2) adding Gaussian noise $\sigma_{k}\epsilon$. For operation (1), we prove with Lemma 7, that the bi-Lipschitz property ensures that applying $\bm{\mu}_{\theta}$ transforms the density bound with updated means and scales the variances. Moreover, the block-preserving property ensures that on-manifold and orthogonal components do not mix, maintaining the factorization. For operation (2), we can observe that adding the aforementioned Gaussian noise simply inflates along and orthogonal manifold components’ variances by $\sigma_{k}^{2}$. Combining (1) and (2) closes the induction and gives, at $t=0$, a block Gaussian lower bound centered at $\mu_{0}$ (Equation 1). We provide the full statement and the formal proof in Appendix B.
Intuitively, this theorem suggests that, the lower bound of the density assigned at a certain point by a well-trained diffusion model depends on how far it is along the manifold from the model’s noise-free center and how far it is off the manifold. Because of the contraction towards the manifold, off-manifold components incurs harsh penalty under the this well-train model, with likelihood dropping rapidly as the off-manifold component grows.
@@ 6 Conditioned Planning with Diffusion
The theoretical analysis in the previous section provides promising signals for OOC generalization in diffusion models. However, it also suggests that reliable generalization requires the linear manifold assumption and a well-trained diffusion model that can capture the characteristics of the manifold. While in environments like a 2D maze, the linear manifold assumption can be naturally satisfied (as 2D planes through the origin are linear subspaces), for more complex environments, this assumption does not necessarily hold true without a powerful encoder that maps the element set to a proper embedding space. Moreover, the block-wise bi-Lipchitz with contraction and block preserving denoising assumption indicates that the diffusion model needs to be manifold (i.e., element combination) aware, which is also not guaranteed with vanilla unconditional diffusion models. As a result, careful model architectural designs are required to ensure empirical performance.
We take inspiration from recent computer vision research, which provides evidence of the combinatorial generalization capabilities of conditioned diffusion models in more complicated data spaces: Aithal et al. (2024) identifies the phenomena where diffusion models generate samples out of the support of training distribution through interpolating different complex modes on a data manifold. Kadkhodaie et al. (2023) demonstrate generalization to unsupported data by showing two diffusion models trained on large enough non-overlapping data converge to the same denoising function. Okawa et al. (2024) showed that given different concepts like shape, color, and size in synthetic shape generation, conditional diffusion models demonstrate a multiplicative emergence of combinatorial abilities where it will first learn how to generalize to concepts closer to the training samples (i.e. only change one of color, shape, and size) and eventually adopt full compositional generalization ability with enough training.
Therefore, we propose to use a conditioned diffusion model with learnable element embeddings to tackle the OOC state problem for decision-making tasks. Following the trajectory-based planning formulation of Janner et al. (2022), we train a conditional diffusion model to denoise (predict) future state-action pairs conditioned on the current observation. Importantly, this training is carried out without reward guidance; instead, it is trained via behavior cloning on successful demonstrations collected by a PPO (Schulman et al., 2017) model in the in-distribution environment, capturing the relationship between states, successful actions, and future states. Conditioning is applied in two complementary ways. First, we enforce consistency of the generated trajectory with the current observation by replacing the corresponding component of the predicted trajectory with the actual environment observation at every denoising step. Second, to enhance this modeling, we incorporate a cross-attention layer after each residual block in the Diffusion UNet, allowing the predicted trajectory to directly attend to the latent vector $\bm{z}$ contained in the observation of the current timestep. Consequently, when conditioned on novel combinations of base elements, the generated trajectories remain dynamically consistent with the underlying dynamics (MDP), facilitating robust generalization to novel scenarios. For deployment, the first action in the generated trajectory is then used to step the environment (Appendix C.4). Detailed architecture of the model can be found in Appendix C.3 and the planning process is described in Algorithm 1.
Note that in practice, our conditioned diffusion model and the element embeddings are trained end-to-end, and thus the resulting learned model does not necessarily fit the theoretical assumptions above. However, as we will show in the next section, empirically this model demonstrates successful OOC generalization in various RL environments.
Input: Diffusion model $\bm{\epsilon_{\theta}}$, compositional elements extractor $r$, learnable embedding function $h$, classifier-free guidance scale $\lambda$, state dimentionality $d_{S}$, initial observation $\bm{o}$, environment simulator envenv
while not done do
  Initialize ${\bm{s}}_{t}\sim\mathcal{N}(0,I)$
  $\bm{c}\leftarrow r(\bm{o})$ $\triangleright$ Extract observed compositional information
  $\bm{z}\leftarrow h(\bm{c})$ $\triangleright$ Obtain element embedding
  for $t\leftarrow T,...1$ do
   ${\bm{s}}_{t}[:d_{S},0]\leftarrow\bm{o}$ $\triangleright$ (Cond state) Replace the first denoised
                              state with observed $\bm{o}$
   $\bm{\widetilde{\epsilon_{t}}}=(1+\lambda)\bm{\epsilon_{\theta}}({\bm{s}}_{t},\bm{z},t)-\lambda\bm{\epsilon_{\theta}}({\bm{s}}_{t},t)$ $\triangleright$ (Cond z) Classifier free guidance
   $\bm{x_{t-1}}=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{s}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\widetilde{\epsilon_{t}}}\right)+\sigma_{t}\bm{\epsilon_{t}},\text{ where }\bm{\epsilon_{t}}\sim\mathcal{N}(0,I)$
  end for
  $\bm{a}\leftarrow\bm{x_{0}}[d_{S}:,0]$ $\triangleright$ Extract action
  $\bm{o}\leftarrow env.step(a)$
end while